OpenStudy (anonymous):

Mandy bought a bag of peanuts to share with her friends. • Trisha received 1/2 of the peanuts in the bag. • Vince received 1/4 of the peanuts Trisha received. • Ray received 1/3 of the peanuts Vince received. If Ray received 4 peanuts, how many peanuts were in the bag Mandy bought? A 13 B 32 C 96 D Not here

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

96 peanuts... 1/2 of 96 = 48 trisha 1/4 of 48 = 12 vince 1/3 of 12 = 4 ray